OpenStudy (anonymous):

The total electric flux E through a sphere is 4.78 *10^7 N m2/C. What is the value of the enclosed charge q?

OpenStudy (anonymous):

\[\oint_S {E_n dA = \frac{Q_{inside}}{{\varepsilon _0 }}} \\ 4.78\,\,{= \frac{Q_{inside}}{{\varepsilon _0 }}} \]

OpenStudy (anonymous):

i cant read that too well im sorry

OpenStudy (anonymous):

the total electric field through surface is\[q/\in \]

OpenStudy (anonymous):

\[ \Huge \oint_S {E_n dA = \frac{Q_{inside}}{{\varepsilon _0 }}} \\ \\ \\ \Huge 4.78\,\,{= \frac{Q_{inside}}{{\varepsilon _0 }}}\]

OpenStudy (anonymous):

so do you use the 4.78? is there a way you can draw it out so i can understand it more?